# GraphGhost: Tracing Structures Behind Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable reasoning capabilities, yet the structural mechanisms underlying these abilities remain under explored. In this work, we introduce GraphGhost, a unified framework that represents neuron activations and their signal propagation as graphs, explaining how LLMs capture structural semantics from sequential inputs and generate outputs through structurally consistent mechanisms. This graph-based perspective enables us to employ graph algorithms such as PageRank to characterize the properties of LLMs, revealing both shared and model-specific reasoning behaviors across diverse datasets. We further identify the activated neurons within GraphGhost and evaluate them through structural interventions, showing that edits to key neuron nodes can trigger reasoning collapse, altering both logical flow and semantic understanding. Together, these contributions position GraphGhost as a powerful tool for analyzing, intervening in, and ultimately understanding the structural foundations of reasoning in LLMs.
@@
Xinnan Dai1, Kai Guo1thanks: Correspondence to guokai1@msu.edu.    The code and data are available at https://github.com/DDigimon/GraphGhost, Chung-Hsiang Lo2, Shenglai Zeng1, Jiayuan Ding4, Dongsheng Luo3,Subhabrata Mukherjee4, Jiliang Tang1 1Michigan State University  2 Northeastern University 3 Florida International University  4 Hippocratic AI {daixinna, guokai1}@msu.edu
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated promising capabilities in capturing structural information from text Sui et al. (2024); Wang et al. (2025, 2023). Notable examples include identifying relationships between entities in context Zhang and Soh (2024), reasoning and planning through structured thinking processes Yao et al. (2023a); Besta et al. (2024); Wu et al. (2024), and recognizing logical connections across paragraphs Pan et al. (2023). Trained on large-scale corpora, recent studies suggest that LLMs can transfer reasoning processes by extracting semantics and generalizing learned reasoning strategies Muennighoff et al. (2025). This raises the question of whether LLMs internally construct reasoning structures and leverage them across diverse tasks.
Figure 1: An illustration of GraphGhost based on Qwen3-0.6B. a) GraphGhost illustrates a structural mechanism within LLMs that operates consistently across datasets. L indicates the layer number. Sequential inputs activate subsets of neurons, which merge tokens in the intermediate layers to capture semantic meaning. These activated neurons then propagate signals to the logit layers to generate outputs. b) Muting the neuron node. Changing the structure of GraphGhost leads to corresponding changes in both language expression and logical reasoning. “等于” means “equal to” in Chinese.
To understand the mechanisms of constructing reasoning structures from text, prior studies have often simplified reasoning tasks into toy examples such as graph-based reasoning to examine how LLMs trace reasoning paths within giving structures Sanford et al. (2024). The underlying hypothesis is that uncovering how decoder-only transformers construct such graphs may yield insights that generalize to real-world tasks Saparov et al. (2025). While these toy graph settings provide preliminary insights—highlighting mechanisms such as token merging Dai et al. (2025b) and edge memorization Wang et al. (2024), they fall short when applied to complex, real-world structured data. To bridge this gap, we introduce GraphGhost, a framework that provides a structural perspective on how LLMs across different architectures perform reasoning tasks in diverse domains. GraphGhost captures neurons and their signal propagation inside LLMs, revealing how neurons are activated across samples and datasets, enabling shared semantic understanding and supporting the construction of reasoning processes across contexts.
As illustrated in Figure 1 a), GraphGhost represents a structural mechanism modeled within LLMs, where neuron activations capture semantic information and guide token outputs. In the sentence-reading stage, LLMs merge tokens to derive semantic meaning, thereby activating specific neurons associated with those tokens. These activated neurons then propagate signals to higher layers, ultimately influencing the logit neurons in the output layers. Notably, while certain neurons are activated for the given samples, many remain silent. These silent neurons are reserved for unseen inputs from different domains, enabling shared output behaviors across contexts. For example, the symbol ‘X’ may activate the ‘=’ symbol in mathematical equations, even if the current sample does not require it, leaving those neurons in a silent state until needed.
We define neurons as nodes and the signal propagation among neurons as edges. This representation allows us to analyze LLMs from a graph perspective. For instance, by inspecting high in-degree nodes or applying PageRank, we can identify key nodes that control next-token prediction, as further elaborated in Section 3. Furthermore, we evaluate the effectiveness of these nodes through structural interventions, such as node removal (Figure 1 b)). Removing specific nodes can lead to semantic shifts, such as switching to another language, or logical changes. In addition, we conduct statistical analyses of these structural perturbations to quantify their effects, as presented in Section 4. In conclusion, our key contributions are summarized as:
  1. 1.
    We introduce GraphGhost, a unified framework represented in graphs, explaining how LLMs capture structural semantics from sequential inputs and generate outputs through a structurally consistent mechanism.
  2. 2.
    We provide a graph-based perspective for understanding LLMs and apply graph algorithms on GraphGhost to characterize the properties of LLMs, revealing both unique and shared behaviors across different models and datasets.
  3. 3.
    We identify the activated neurons within LLMs and evaluate them by editing the structure of GraphGhost. Our experiments show that key neuron nodes can trigger reasoning collapse, altering both the logical flow and semantic understanding.
2 Related work and Preliminary
GraphGhost is built on circuit tracing, which is a typical LLM interpretation tool. In this section, we first introduce the background on LLM interpretability and reasoning tasks in Section 2.1 and Section 2.2, respectively. To illustrate how LLMs construct internal structures that align with explicit reasoning texts, we design a toy experiment in Section 2.3 to illustrate the alignment of inner signal propagation with explicit structure descriptions. Building on this foundation, we provide the preliminary definition in Section 2.4.
2.1 Interpretation of LLMs
Interpreting how LLMs perform reasoning is essential for aligning their behavior with human understanding Zhao et al. (2024). Existing approaches fall into two categories, self-explanation Gu et al. (2024) and causal-mechanistic interpretation Bloom et al. (2024). The details are suggested in Appendix A.1
The causal-mechanism approach focuses on explaining how each token is predicted Anthropic (2023); Ameisen et al. (2025). Based on learning the internal reasoning pathways, Lindsey et al. (2025) introduced attribution graphs to capture token flows that contribute to final predictions. Formally, an attribution graph is defined as $G_{A}=(\hat{V},\hat{E},A)$, where $\hat{V}$ denotes activated neurons, $A_{s\to t},;s,t\in\hat{V}$ represents the contribution from node $s$ to $t$ (computed via backward Jacobians), and $\hat{E}$ indicates whether an edge exists. The edge set is controlled by an edge ratio, which filters the top influential connections according to $A$.
2.2 Reasoning Tasks
Reasoning tasks including mathmatical Cobbe et al. (2021), scientific Khot et al. (2020), and logic reasoning Clark et al. (2019) are set to evaluate the ability in solving complex problems. The details are in Appendix C.2
Graph reasoning represents a distinct type of reasoning task Dai et al. (2025a). In this setting, models are provided with textual descriptions of graph structures and asked to reason over them. Success on such tasks suggests that LLMs can leverage implicit structural representations to guide reasoning Wang et al. (2023). However, how these implicit structures are internally constructed and maintained within LLMs remains an open question.
2.3 Path tracing in the decoder-only transformer
Graph reasoning tasks, particularly path reasoning, offer a prototypical setting for understanding how structural information in sequential inputs can be mapped to outputs that describe those structures Wang et al. (2023). In this task, the model is provided with a graph description and a query consisting of a start node and an end node, and the objective for LLMs or decoder-only transformers is to predict the path connecting the two nodes Dai et al. (2024); Wang et al. (2024). We adopt this basic setup to investigate how decoder-only transformers perform path reasoning. Following prior work Wang et al. (2024), we use simplified graph descriptions, and a query (“from start point S to end point E”) and must predict the connecting path. as illustrated in Figure 2(a), and train a 5-layer GPT-2 architecture decoder-only transformer on this task. The trained model achieves 98% accuracy on the test set. We then apply the circuit-tracing method, specifically using transcoder Dunefsky et al. (2024), to analyze how the transformer models the path-finding process, as shown in Figure 2(b). Full experimental details are provided in Appendix C.1.
@@ 2.3 Path tracing in the decoder-only transformer
Circuit tracing reveals how decoder-only transformers review the previous context to identify related nodes and predict the next token. From a structural understanding perspective, neighboring nodes are merged in the middle layers. For example, at the third layer, node 4 is merged with node 5, forming the semantic representation of the edge between them. Further, in generating the output tokens, the final token in the sequence receives the message that an edge exists between 4 and 5 at layer 3, leading the model to predict 5 as the next token.
From this toy example, we observe that decoder-only transformers are not only capable of reasoning about paths within given graph structures, but that their underlying mechanism involves activated neurons propagating signals across layers. However, beyond the toy example with decoder-only transformers, does this mechanism extend to LLMs that can solve more complex tasks expressed in natural language? To address this, we introduce GraphGhost, a unified framework that represents these neuron activations and their signal flows as graphs. GraphGhost enables us to trace and mergeattribution graphs defined in Section 2.1 across diverse reasoning tasks, revealing how LLMs encode and reuse structural reasoning patterns. In the following sections, we apply GraphGhost to multiple models and domains to analyze how their internal reasoning structures build up, interact, and generalize.
@@
2.4 Preliminary definitions
We formalize the core concepts underlying GraphGhost. GraphGhost is constructed from attribution graphs obtained using $\text{CircuitTracer}(\cdot)$, which traces neuron activations contributing to next-token prediction and records their signal flow. Let $X=(t_{1},\dots,t_{n})$ be the input tokens of an LLM with layers $\{f^{(1)},\dots,f^{(L)}\}$. For a target token $t_{n+1}$, the attribution graph is defined as $A_{n+1}=\text{CircuitTracer}(t_{n+1}),$ where each node represents an activated neuron $(t_{i},\ell)$ with activation $h^{(\ell)}(t_{i})=f^{(\ell)}(h^{(\ell-1)}(t_{i}))$, and each directed edge $w_{s\to t}$ encodes the contribution from neuron $s$ to $t$. The set of contributing tokens is $\hat{T}(t_{n+1})=\{\,t_{i}\mid(t_{i},\ell)\in A_{n+1}\,\}$. where $\ell$ denotes the layer index. These local attribution graphs form the building blocks of GraphGhost, which aggregates them across samples into a unified weighted structure.
Following the definition of attribution graphs in Ameisen et al. (2025), a neuron node is represented as a pair $(t_{i},\ell)$, where $t_{i}$ denotes a token and $\ell$ denotes the layer index. Each neuron node has a corresponding hidden activation $h^{(\ell)}(t_{i})$, representing the hidden state of token $t_{i}$ at layer $\ell$, computed as $h^{(\ell)}(t_{i})=f^{(\ell)}\big(h^{(\ell-1)}(t_{i})\big).$ For a source neuron node $s=(t_{i},\ell_{s})$ and a target neuron node $t=(t_{j},\ell_{t})$ with $\ell_{s}<\ell_{t}$, the edge weight in the attribution graph is defined as $w_{s\to t}=A_{s\to t},$ where $A_{s\to t}$ denotes the attribution score from $s$ to $t$.
For example, in Figure 2,to generate token $5$, the contributing tokens are $\{4,5,7,\texttt{"|"},\texttt{"S"},\texttt{"P"}\}$. The corresponding neuron nodes include $h^{(3)}(5)$, $h^{(4)}(7)$, and $h^{(5)}(4)$, among others. The edge weight from $h^{(3)}(5)$ to $h^{(5)}(4)$ is denoted as $w_{h^{(3)}(5)\to h^{(5)}(4)}$.
3 GraphGhost in LLMs
Beyond the toy examples, we apply GraphGhost to real-world LLMs to uncover their internal reasoning structures. GraphGhost traces neuron activations that contribute to next-token generation, represents them as local attribution graphs, and merges these across samples into a global weighted graph. This unified structure reveals stable signal propagation patterns shared across tasks and models. Section 3.1 details the tracing and aggregation process. To evaluate whether these structures correspond to meaningful reasoning behaviors, we introduce the experimental setup in Section 3.2. which specifies the models, datasets, and tracing configurations used for constructing GraphGhost. Section 3.2 and Section 3.4 then present case studies and quantitative analyses based on these settings.
3.1 Tracing the GraphGhost
To construct a unified GraphGhost that captures reasoning structures shared across multiple samples, we aggregate the attribution graphs obtained from each sample. Given a set of samples $\mathcal{S}=\{s_{1},s_{2},\dots,s_{m}\}$, where each sample $s_{i}$ contains an answer token $t_{n}^{(i)}$, we initialize a global graph $G=(V,E,W)$ with empty node, edge, and weight sets. For each sample $s_{i}$, we start from its answer token $t_{n}^{(i)}$ and iteratively trace the contributing tokens and their local attribution graph using the $\text{CircuitTracer}(\cdot)$ function. Each traced graph produces a set of neuron nodes $\hat{V}$ and directed edges $\hat{E}$, which are merged into the global GraphGhost. If an edge $e\in\hat{E}$ does not exist in $E$, it is added with an initial weight $W(e)=1$; otherwise, its weight is incremented by one. This cumulative weighting reflects the frequency of edge occurrences across samples, capturing the stability and recurrence of signal propagation pathways. After processing all samples, the resulting GraphGhost $G=(V,E,W)$ represents a multi-sample structural abstraction of the model’s internal reasoning dynamics, which is a weighted directed graph describing neuron-level signal propagation. The full procedure is summarized in Algorithm 1 fromAppendix B.
3.2 Experiments setting
We evaluate a range of LLMs and datasets to investigate whether GraphGhost constitutes a general mechanism across models. Specifically, we consider Qwen3-0.6B (28 layers), Qwen2.5-1.5B (28 layers), Llama3.1-8B-Instruct (32 layers), DeepSeek-distilled-Qwen2.5B-1.5B (DS-Qwen-1.5B) (28 layers), and DeepSeek-distilled-Llama-8B (DS-Llama-8B) (32 layers). For datasets, we examine three domains: (i) mathematical reasoning (GSM8K, MAWPS), (ii) logical reasoning (ProntoQA, BoolQA), and (iii) scientific reasoning (ARC-Easy, QASC). To elicit the models’ reasoning processes, we apply Chain-of-Thought prompting. Following Hanna et al. (2025), we then train transcoders at every layer for each dataset–model pair to probe neuron activations during reasoning, ensuring that the transcoders effectively capture the most strongly activated neurons. All the experiments are run on a single H200 GPU. Full experimental details are provided in Appendix C.2.
3.3 Case study on GraphGhost
@@
Figure 3: Case study of tracing contributing tokens for the answer 14 in Qwen3-0.6B model. We highlight the numerical reasoning flow captured by GraphGhost to visualize how number-related tokens contribute to the final prediction.
@@
We further provide some case studies of GraphGhost on contributing nodes and substructures, respectively. Figure 3 is an illustrative example of tracing contributing tokens within a single sample. The identified token set $T$ includes tokens from both the input question (highlighted in green) and those generated along the reasoning chain (in orange). As shown in the figure, number-related tokens emerge through three distinct mechanisms: (i) step-by-step progression (e.g., 1–2–3–4), (ii) direct copying from the input context (e.g., 30 and 28), and (iii) generation influenced by neighboring tokens, such as the digit “1” in the final answer “14”, which is activated by “=". This tracing process demonstrates how local dependencies between tokens collectively form reasoning paths during sequential generation.
Figure 4 demonstrates a representative subgraph extracted from the GraphGhost constructed on Qwen3-0.6B model, integrating neuron activations from mathematical reasoning (GSM8K and MAWPS), scientific reasoning (ARC), and logical reasoning (ProntoQA). Each neuron node is annotated with its corresponding token and layer ID. For example, the activated period at layer 22 is denoted as $_{22}.$ This figure highlights how neuron activations from diverse domains converge into a unified structural representation, revealing consistent signal propagation patterns underlying different types of reasoning tasks.

### v2

@@ Abstract
Large Language Models (LLMs) exhibit strong reasoning capabilities on structured tasks, yet the internal mechanisms underlying such behaviors remain poorly understood. Existing interpretation methods mainly focus on token-level attributions, which provide limited insight into multi-step reasoning inside the model. We propose GraphGhost, a graph-based framework that models internal token interactions and neuron activations in LLMs as graphs. By aggregating token dependencies traced across layers, GraphGhost captures global information flow underlying model predictions. We formalize GraphGhost from two complementary perspectives: a sample view, which traces token dependencies for individual predictions, and a dataset view, which aggregates recurring structural patterns learned during training. Through graph analytics and quantitative experiments, we show that graph structural properties are closely associated with influential tokens and neuron nodes, and that perturbations to structurally critical nodes lead to measurable changes in reasoning behavior. These results indicate that the structural patterns captured by GraphGhost reflect meaningful internal organization of LLM reasoning. The codes are available at software part. Artifacts will be made available for research use only.
@@
Xinnan Dai1, XianXuan Long1, Chung-Hsiang Lo2, Kai Guo1thanks: Correspondence to guokai1@msu.edu.    The code and data are available at https://github.com/DDigimon/GraphGhost, Shenglai Zeng1, Dongsheng Luo3, Jiliang Tang1 1Michigan State University  2 Northeastern University 3 Florida International University {daixinna, guokai1}@msu.edu
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in solving structured reasoning tasks, where solutions require the composition of multiple intermediate steps governed by explicit relational or logical dependencies, such as mathematical problem solving, logical inference, and graph-based reasoning (Bommasani et al., 2022; Pan et al., 2023). Despite this success, the internal mechanisms by which LLMs represent, compose, and reuse such structural information during reasoning remain poorly understood (Nanda et al., 2023; Nananukul et al., 2025; Pandey et al., 2025). This is important because it provides a principled way to understand the source of intelligence in LLMs.
Figure 1: From local explanations and human-aligned reasoning to global internal structures. (a) Token-level attribution explanations provide localized views of individual predictions. (b) Human-aligned step-by-step reasoning produces coherent reasoning traces without revealing internal mechanisms. (c) GraphGhost captures global token interactions inside LLMs as a structured graph, enabling mechanistic analysis of internal reasoning processes.
Most existing efforts to understand LLMs’ behaviors on structured reasoning tasks focus on token-level explanations (Chen et al., 2025; Tak et al., 2025), as illustrated in Figure 1(a), which trace individual predictions back to contributing tokens or activations and provide only localized views of isolated outputs. Given that reasoning capability plays a central role in LLM performance across reasoning tasks (Wei et al., 2022; Lewkowycz et al., 2022) as shown in Figure 1 (b), existing token-level approaches provide limited insights into how representations are composed and reused across multiple reasoning steps (Geiger et al., 2025; Dziri et al., 2023). This limitation is increasingly recognized as a major obstacle to meaningful interpretability, as faithful understanding requires alignment between internal reasoning mechanisms and human-interpretable reasoning logic (Zhao et al., 2024; Turpin et al., 2023; Barez et al., 2025).
To bridge this gap, we introduce GraphGhost, a graph-based framework that captures internal token interactions and neuron activations in LLMs as graphs as shown in Figure 1(c). Motivated by the central role of step-by-step reasoning in LLM performance, GraphGhost moves beyond isolated token attributions to explore how the reasoning steps are driven under the individual token explanation. This structural representation enables us to analyze reasoning as an internal, mechanistic process, revealing how intermediate tokens are merged, propagated, and reused during generation.
Specifically, in Section 3, we formalize GraphGhost from two complementary perspectives. First, in the sample view, GraphGhost traces internal token dependencies within individual input–output pairs that contribute to a specific prediction, exposing reasoning structures for LLMs. Second, in the dataset view, GraphGhost is constructed by aggregating structures across samples to reveal recurring and shared token-level patterns learned from data. Through graph analytics, we demonstrate that LLM reasoning behavior is tightly coupled to internal graph structural properties, with influential tokens and neuron nodes characterized by structurally critical roles. These findings suggest that internal graph structures fundamentally shape how individual tokens influence LLM reasoning, through their structural roles within the model.
In Section 4, we conduct extensive quantitative experiments to assess whether the structural patterns captured by internal graphs are meaningfully tied to LLM reasoning behavior. Using fidelity, infidelity, and sparsity metrics, we demonstrate that the tokens and neuron nodes identified by GraphGhost exert a significantly strong influence on model behavior. Furthermore, targeted interventions on structurally important nodes lead to measurable changes in reasoning trajectories and outputs, providing empirical evidence that GraphGhost captures meaningful internal structures governing LLM reasoning.
In conclusion, we summarize our key contributions as follows:
1) We propose GraphGhost, a unified framework for understanding how LLMs perform reasoning by modeling internal token interactions and neuron activations at both the sample and dataset levels.
2) Using GraphGhost, we reveal that internal graphs capture structural constraints that govern model behavior, shaping information composition and propagation during reasoning.
3) We conduct quantitative experiments to test whether graphs constructed by GraphGhost capture meaningful structural patterns of LLM reasoning. Our results show that structurally important token-related neuron nodes strongly influence reasoning trajectories, and that intervening on these nodes alters both language expression and logical flow.
2 Related Work
While Chain-of-Thought (CoT) prompting enhances LLM reasoning (Wei et al., 2022; Kojima et al., 2022), its faithfulness to the model’s intrinsic computations remains contested (Jacovi and Goldberg, 2020; Lyu et al., 2023). Prior work shows that CoT often acts as post-hoc rationalization, generating explanations that diverges from the underlying causal logic (Turpin et al., 2023) or internal attention patterns (Lanham et al., 2023; Barez et al., 2025). Moreover, such behavioral analyses provide limited insight into the specific neurons or circuits driving model predictions.
To overcome the lack of causal rigor in probing Alain and Bengio (2018); Belinkov (2021) and attention analysis Clark et al. (2019); Vig and Belinkov (2019), research has pivoted to attribution methods (Sundararajan et al., 2017; Meng et al., 2023) and circuit tracing (Elhage et al., 2021). Recent methods leverages Sparse Autoencoders (SAEs) Bricken et al. (2023); Cunningham et al. (2023) and Transcoders Dunefsky et al. (2024) to decompose dense representations into interpretable features (Bloom et al., 2024; Ferrando et al., 2025; Paulo et al., 2025), enabling attributed graphs (Hanna et al., 2025; Marks et al., 2025) to model causal flow. GraphGhost extends this framework to uncover internal structures that affect LLM behavior in both sample and dataset view.
Figure 2: We use a simple example: 2 + 2 * 4. (a) An attributed graph interpreting a single-token prediction. (b) GraphGhost (sample view) aggregates token dependencies along the reasoning chain. (c) GraphGhost (dataset view) captures cross-layer token correlations learned from the corpus. Here, “2_1” denotes token 2 at layer 1; “-_2” does not appear in the sample but is activated from other corpus contexts. Token “2_1” participates in both “2*4” and “2+8” computations.
3 Tracing the GraphGhosts
The reasoning behavior of LLMs emerges through successive token generations. Accordingly, existing interpretation methods primarily operate at the token level, attributing individual predictions to specific tokens or activations and providing localized explanations, as illustrated in Figure 2(a) Ameisen et al. (2025). However, reasoning does not arise from isolated tokens alone, but from structured interactions among multiple tokens within a sample and from recurring patterns shared across samples.
To capture these aspects, we introduce two complementary views: a sample view and a dataset view. The sample view models token interactions within a single reasoning instance and formalizes this instance-specific structure as a GraphGhost (Figure 2(b); Section 3.1). The dataset view aggregates such structures across samples to reveal reusable reasoning patterns and is formalized in Section 3.2 (Figure 2(c)).
3.1 Sample View
LLMs have demonstrated strong capabilities in chain-of-thought reasoning. However, whether these reasoning chains faithfully correspond to coherent internal reasoning mechanisms remains unclear. Rather than attempting to answer this question conclusively, we begin with a toy example to illustrate how circuit-tracing methods can be used to probe the internal computations underlying such reasoning behaviors.
Toy Example on Graph Reasoning
Since search paths on a graph naturally decompose reasoning into node-by-node transitions, we adopt the path-reasoning task as a toy example to illustrate how nodes are selected at each step of the reasoning process. Following prior work Saparov et al. (2025); Wang et al. (2024) that studies LLM reasoning and planning through simplified graph abstractions, we focus on shortest-path search as our canonical setting, as illustrated in Figure 3(a).
Graphs are specified as lists of edges, with the symbol “|” used to separate individual edges. In the queries, we denote the start and end nodes using the markers “S” and “E”, respectively. In the answers, we use the symbol “P” to indicate that the model is required to generate an explicit reasoning path. we train a 5-layer GPT-2 architecture decoder-only transformer on this task. The trained model achieves 98% accuracy on the test set.To understand how the trained model internally performs path finding, we analyze its inference process using circuit-tracing methods.
Specifically, following Hanna et al. (2025), we train a transcoder Dunefsky et al. (2024) for each layer of the transformer. Each transcoder takes the layer-wise hidden activations as input and learns a sparse feature representation that reconstructs these activations, yielding token-level semantic features for each MLP layer. We then replace the original MLP layers with the corresponding transcoders and compute the logit-level contribution weights of each transcoder feature to the final prediction. Using transcoder features as nodes and their logit contribution weights as directed edges, we construct the attributed graph. The resulting attribution graph, visualized in Figure 3(b), illustrates how the transformer models the path-finding process. Full experimental details are provided in Appendix B.1.
@@ 3.1 Sample View
Circuit tracing reveals that decoder-only transformers revisit the preceding context to identify relevant nodes and predict the next token. From a structural perspective, neighboring nodes are progressively merged in the middle layers. For example, at the third layer, node 4 is merged with node 5, forming a semantic representation of the edge between them. During generation, the final token in the sequence receives the signal that an edge exists between nodes 4 and 5 at layer 3, which leads the model to predict node 5 as the next token. This mechanism differs from human graph reasoning, which explicitly retrieves relations, checks connectivity, and composes shortest paths. We then evaluate this assumption on real LLM reasoning tasks.
@@
GraphGhost in Sample Interpretation
In the toy example, we observe that reasoning in basic decoder-only transformers is driven by token-level associations, where previously observed in-context tokens contribute to information aggregation during token generation, rather than through explicitly modeled logical structures. Motivated by this observation, we merge answer-related circuits to examine how reasoning is internally realized for a specific query in LLMs.
At the sample level, GraphGhost aims to capture the internal structural dependencies that contribute to a model’s final prediction. Given a sample $s$ with final answer token $t_{n}$, we initialize an empty graph $G^{\text{sample}}=(V,E)$ together with a local edge-weight accumulator $W^{\text{local}}$, which is used to aggregate contribution strengths when the same dependency is discovered multiple times during backward expansion. The graph construction starts from $t_{n}$ and iteratively expands backward through the model’s internal computation.
For each token $t$ in the current frontier that is not part of the question tokens, we apply a circuit tracing procedure to extract the local logit-based attribution graph rooted at $t$ follows the previous paragraph. This step yields a set of contributing tokens as sample-view nodes $\hat{V}$, directed edges $\hat{E}$, and local attribution weights $W^{\text{loc}}$, characterizing how information flows to the prediction of $t$. Newly discovered tokens are added to the frontier to enable recursive exploration of upstream dependencies.
The nodes and edges from each local attribution graph are by taking the union over nodes and edges, removing duplicates and retaining only the existence of dependency relations. As a result, the sample-level GraphGhost is an unweighted structural graph that captures which token-level dependencies are present, independent of their individual contribution strengths. In parallel, local edge weights are accumulated in $W^{\text{local}}$ for each observed edge, which is retained solely for later aggregation at the dataset level. The final output is a sample-specific GraphGhost graph $G^{\text{sample}}=(V,E)$ together with the local edge-weight map $W^{\text{local}}$. The overall algorithm is shown in Algorithm 1 in Appendix A.
@@
Figure 4: Case study tracing tokens contributing to the answer 14 in Qwen3-0.6B, illustrating how GraphGhost captures the numerical reasoning flow linking number-related tokens to the final prediction.
@@
The sample view of GraphGhost characterizes step-by-step reasoning in LLMs by modeling structural dependencies among tokens as reasoning unfolds. Here, we show an example in Figure 4. The identified token set $T$ includes tokens from both the input question (highlighted in green) and those generated along the reasoning chain (in orange). As shown in the figure, number-related tokens emerge through three distinct mechanisms: (i) step-by-step progression (e.g., 1–2–3–4), (ii) direct copying from the input context (e.g., 30 and 28), and (iii) generation influenced by neighboring tokens, such as the digit “1” in the final answer “14”, which is activated by “=". This tracing process demonstrates how local dependencies between tokens collectively form reasoning paths during sequential generation.
3.2 Dataset View
The reasoning capabilities of LLMs are learned from large-scale corpora, which motivates the assumption that certain neuron units within LLMs encode reusable token-level connections that support generalization to new reasoning tasks. Accordingly, we introduce the dataset-level view of GraphGhost to explore the implicit structures inside LLMs.
3.2.1 Method
To build the dataset-level of GraphGhost, we aggregate sample-level reasoning structures to reveal recurring internal information-flow patterns shared across multiple samples from datasets. Given a set of samples $\mathcal{S}=\{s_{1},\ldots,s_{m}\}$, we construct a unified graph for the dataset as $G^{\text{data}}=(V^{\text{data}},E^{\text{data}},W^{\text{data}})$.
For each sample $s_{i}\in\mathcal{S}$, we first obtain its sample-level GraphGhost $G^{(i)}=(V^{(i)},E^{(i)})$ together with a local edge-weight map $W^{\text{local}}$ using the sample-view construction. The dataset-level node and edge sets are formed by merging sample graphs via set union: $V^{\text{data}}=\bigcup_{i}V^{(i)}$ and $E^{\text{data}}=\bigcup_{i}E^{(i)}$. In parallel, edge attribution strengths are accumulated across samples. For each edge $e\in E^{(i)}$, we update its dataset-level weight as $W^{\text{data}}(e)\leftarrow W^{\text{data}}(e)+W^{\text{local}}(e),$ where $W^{\text{local}}$ captures unnormalized, sample-specific contribution strengths. This accumulation preserves both the frequency and strength of recurring dependencies across samples. After aggregating all samples, we apply row-stochastic normalization to $W^{\text{data}}$ to obtain relative edge importance at the dataset level. The resulting graph captures recurring reasoning paths, while the normalized weights quantify their global importance across the dataset. The full procedure is summarized in Algorithm 2 in Appendix A.
The dataset view aims to reveal whether LLMs exhibit recurring internal structures across different tasks and datasets. To this end, we visualize a subgraph extracted from the dataset-level GraphGhost in Figure 5, which aggregates sample-level GraphGhosts constructed from four datasets spanning different domains. In this graph, each node represents a token–layer pair identified by GraphGhost, and node colors indicate the datasets in which the token participates, while mixed colors denote tokens shared across multiple datasets. From this visualization, several consistent patterns emerge. First, certain tokens appear across all four datasets, suggesting shared internal mechanisms. For example, the period symbol “.” at layer 22 exhibits strong logit contributions across all datasets, and the discourse marker “So” frequently appears in diverse reasoning chains. Second, domain-specific patterns are also evident: arithmetic-related tokens such as “Adding” and expressions like “6+” are shared between GSM8K and MAWPS, which both involve numerical reasoning. In contrast, some nodes are task-specific; for example, $\mathrm{B_{18}}$ appears only in ARC, reflecting choice-based reasoning. This shows that dataset-level GraphGhost captures both shared reasoning primitives and task-specific structures.